Estimate the value of the equilibrium constant at 655 K for each of the following reactions. ΔG∘f for BrCl(g) is −1.0 kJ/mol. The standard molar entropy, S∘, for BrCl(g) is 240.0 J/mol⋅K.

PART A: 2NO2(g)⇌N2O4(g). ΔH∘f for N2O4(g) is 9.16 kJ/mol.

PART B: Br2(g)+Cl2(g)⇌2BrCl(g) .ΔH∘f for BrCl(g) is 14.6 kJ/mol.
Answer #1

Part-A

2NO2(g) --------------- N2O4(g)

ΔH∘ rxn = ΔH0f of N2O4 - 2 x ΔH∘ f of NO2

ΔH∘ rxn = 9.16 - 2 x33.18

ΔH∘rxn = - 57.2KJ

ΔS0 rxn = ΔS0f of N2O4 - 2 xΔS0f of NO2

ΔS0rxn = 304.18 - 2x 239.95

ΔS0 rxn = - 175.72J

ΔG0rxn = ΔH0rxn - T ΔS0 rxn

ΔG0rxn = - 57.2 KJ - 298x(-175.72)

ΔG0rxn = - 4835.44 K

ΔG0rxn = - 4.835 KJ

ΔG0rxn = - RTlnK                 R= 8.314x10^-3KJ             T= 655K

- 4.835 = - 8.314x10^-3 x 655x lnK

lnK = 0.888

K = e^0.889

K = 2.43.

part-B

Br2(g) + Cl2(g) --------------- 2 BrCl(g)

ΔG0rxn = -1.0KJ

ΔG0rxn = - RTlnK

-1.0 = - 8.314x10^-3 x 655x lnK

lnK = 0.183

K = e^0.183

K= 1.20